OpenStudy (anonymous):

Write the number of inches needed to represent 500 miles if 1/4 in. = 25 miles.

OpenStudy (anonymous):

We have 1/4 in. = 25 miles Therefore, 1 in = 100 miles Or, 100 miles = 1in Therefore, 500 miles = 5* 100 miles = 5 * 1 in = 5 in

OpenStudy (anonymous):

Thank you. :)

OpenStudy (anonymous):

You're welcome. :)